# Teleportation Fidelity of Binary Tree Quantum Repeater Networks

### v2

@@ A.2 Undirected Binary Trees
$l_{d}^{(r)}=\begin{cases}2d,&\text{if}\leavevmode\nobreak\ r=1\\ 1+3(d-1),&\text{if}\leavevmode\nobreak\ r=2\\ 2+4(d-r+1),&\text{if}\leavevmode\nobreak\ 3\leq r\leq d+1\\ 0,&\text{otherwise}.\end{cases}$ (27)
@@
$\left.F^{\text{tel}}_{\text{avg}}(N,p)\right|_{\text{UABT}}$ $=$ $\frac{1}{2}+\frac{2p(N-1)-(N+3)p^{2}-4p^{3}-(N-1)p^{4}+4p^{\frac{N+3}{2}}+4p^{\frac{N+5}{2}}}{2N(N-1)(1-p)^{2}}.$ (30)
@@
Here again, $h_{r}$ denotes the number of $r$-length paths starting from the base vertex 1, and at each level, there are two choices for selecting a vertex to include in the path up to depth $r$. So, $h_{r}=2^{r}$. Now, let us denote by $l_{d}^{(r)}$ the number of $r$-length paths in a USBT with depth $d$. In this case, for $r>2d$, there is no $r$-length path as it is undirected and symmetric tree, and hence $l_{d}^{(r)}=0$ for $r>2d$.
@@
$l_{d}^{(r)}=\begin{cases}3\cdot 2^{d+(\frac{r}{2})-1}-(2d-r+5)\cdot 2^{r-2},&\text{if}\leavevmode\nobreak\ d<r\leq 2d\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ r\leavevmode\nobreak\ \text{is even}\\ 2^{d+(\frac{r+1}{2})}-(2d-r+5)\cdot 2^{r-2},&\text{if}\leavevmode\nobreak\ d<r\leq 2d\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ r\leavevmode\nobreak\ \text{is odd}\\ 3\cdot 2^{d+(\frac{r}{2})-1}-(r+3)\cdot 2^{r-2},&\text{if}\leavevmode\nobreak\ r\leq d\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ r\leavevmode\nobreak\ \text{is even}\\ 2^{d+(\frac{r+1}{2})}-(r+3)\cdot 2^{r-2},&\text{if}\leavevmode\nobreak\ r\leq d\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ r\leavevmode\nobreak\ \text{is odd}\\ 0,&\text{otherwise}.\end{cases}$ (40)
@@
$\left.F^{\text{tel}}_{\text{avg}}(d,p)\right|_{\text{USBT}}$ $=\frac{\left[\begin{aligned} &\sum\limits_{\begin{subarray}{c}r\leq d\\ r\leavevmode\nobreak\ \text{even}\end{subarray}}\left(3\cdot 2^{d+\frac{r}{2}-1}-(r+3)\cdot 2^{r-2}\right)\cdot\left(\frac{1+p^{r}}{2}\right)+\sum\limits_{\begin{subarray}{c}r\leq d\\ r\leavevmode\nobreak\ \text{odd}\end{subarray}}\left(2^{d+\frac{r+1}{2}}-(r+3)\cdot 2^{r-2}\right)\cdot\left(\frac{1+p^{r}}{2}\right)+\\ &\sum\limits_{\begin{subarray}{c}d<r\leq 2d\\ r\leavevmode\nobreak\ \text{even}\end{subarray}}\left(3\cdot 2^{d+\frac{r}{2}-1}-(2d-r+5)\cdot 2^{r-2}\right)\cdot\left(\frac{1+p^{r}}{2}\right)+\sum\limits_{\begin{subarray}{c}d<r\leq 2d\\ r\leavevmode\nobreak\ \text{odd}\end{subarray}}\left(2^{d+\frac{r+1}{2}}-(2d-r+5)\cdot 2^{r-2}\right)\cdot\left(\frac{1+p^{r}}{2}\right)\end{aligned}\right]}{\left[\begin{aligned} &\sum\limits_{\begin{subarray}{c}r\leq d\\ r\leavevmode\nobreak\ \text{even}\end{subarray}}\left(3\cdot 2^{d+\frac{r}{2}-1}-(r+3)\cdot 2^{r-2}\right)+\sum\limits_{\begin{subarray}{c}r\leq d\\ r\leavevmode\nobreak\ \text{odd}\end{subarray}}\left(2^{d+\frac{r+1}{2}}-(r+3)\cdot 2^{r-2}\right)+\\ &\sum\limits_{\begin{subarray}{c}d<r\leq 2d\\ r\leavevmode\nobreak\ \text{even}\end{subarray}}\left(3\cdot 2^{d+\frac{r}{2}-1}-(2d-r+5)\cdot 2^{r-2}\right)+\sum\limits_{\begin{subarray}{c}d<r\leq 2d\\ r\leavevmode\nobreak\ \text{odd}\end{subarray}}\left(2^{d+\frac{r+1}{2}}-(2d-r+5)\cdot 2^{r-2}\right)\end{aligned}\right]}$
$=\frac{1}{2}+\frac{1}{2}\frac{\sum\limits_{\begin{subarray}{c}r=1\\ r\leavevmode\nobreak\ \text{even}\end{subarray}}^{2d}3\cdot 2^{d+\frac{r}{2}-1}\cdot p^{r}+\sum\limits_{\begin{subarray}{c}r=1\\ r\leavevmode\nobreak\ \text{odd}\end{subarray}}^{2d}2^{d+\frac{r+1}{2}}\cdot p^{r}-\sum\limits_{r=1}^{d}(r+3)\cdot 2^{r-2}\cdot p^{r}-\sum\limits_{r=d+1}^{2d}(2d-r+5)\cdot 2^{r-2}\cdot p^{r}}{\sum\limits_{\begin{subarray}{c}r=1\\ r\leavevmode\nobreak\ \text{even}\end{subarray}}^{2d}3\cdot 2^{d+\frac{r}{2}-1}+\sum\limits_{\begin{subarray}{c}r=1\\ r\leavevmode\nobreak\ \text{odd}\end{subarray}}^{2d}2^{d+\frac{r+1}{2}}-\sum\limits_{r=1}^{d}(r+3)\cdot 2^{r-2}-\sum\limits_{r=d+1}^{2d}(2d-r+5)\cdot 2^{r-2}}$
$=\frac{1}{2}+\frac{1}{2}\frac{\sum\limits_{k=1}^{d}3\cdot 2^{d+k-1}\cdot p^{2k}+\sum\limits_{k=1}^{d}2^{d+k}\cdot p^{2k-1}-\sum\limits_{r=1}^{d}(r+3)\cdot 2^{r-2}\cdot p^{r}-\sum\limits_{t=1}^{d}(d-t+5)\cdot 2^{d+t-2}\cdot p^{d+t}}{\sum\limits_{k=1}^{d}3\cdot 2^{d+k-1}+\sum\limits_{k=1}^{d}2^{d+k}-\sum\limits_{r=1}^{d}(r+3)\cdot 2^{r-2}-\sum\limits_{t=1}^{d}(d-t+5)\cdot 2^{d+t-2}}$
$=\frac{1}{2}+\frac{1}{2}\frac{\left(3\cdot 2^{d-1}+\frac{2^{d}}{p}\right)\sum\limits_{k=1}^{d}(2p^{2})^{k}-\left(\frac{3}{4}+(d+5)\cdot 2^{d-2}\cdot p^{d}\right)\sum\limits_{k=1}^{d}(2p)^{k}+\left(2^{d-2}\cdot p^{d}-\frac{1}{4}\right)\sum\limits_{k=1}^{d}k(2p)^{k}}{\left(3\cdot 2^{d-1}+2^{d}\right)\sum\limits_{k=1}^{d}2^{k}-\left(\frac{3}{4}+(d+5)\cdot 2^{d-2}\right)\sum\limits_{k=1}^{d}2^{k}+\left(2^{d-2}-\frac{1}{4}\right)\sum\limits_{k=1}^{d}k2^{k}}$
$=\begin{cases}\displaystyle\frac{1}{2}+\frac{14(2^{d}-1)-d(d+8)}{8(2^{d}-1)(2^{d+1}-1)},&\text{if }\leavevmode\nobreak\ p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{d\cdot 2^{d+1}-\sqrt{2}(2^{d/2}-1)\left[(\sqrt{2}-1)\left\{3+5\cdot 2^{d/2}\right\}+(2^{d/2}-1)\right]}{8(3-2\sqrt{2})(2^{d}-1)(2^{d+1}-1)},&\text{if }\leavevmode\nobreak\ p=\frac{1}{\sqrt{2}},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{2^{d-1}p(3p+2)\left((2p^{2})^{d}-1\right)}{(2p^{2}-1)(2^{d}-1)(2^{d+1}-1)}-\frac{p\left((2p)^{d}-1\right)\left[(2p-1)\left\{3+5\cdot(2p)^{d}\right\}+\left((2p)^{d}-1\right)\right]}{4(2p-1)^{2}(2^{d}-1)(2^{d+1}-1)},&\text{otherwise},\end{cases}$ (41)
$=\frac{1}{2}+\varepsilon(d).$ (42)
@@ Appendix B Large $N$ limit
To understand the large $N$ behaviour for all the cases of binary trees, we consider the formula of $F_{\mathrm{avg}}^{\mathrm{tel}}$. For any value of $p$, we can write the formula of $F_{\mathrm{avg}}^{\mathrm{tel}}(d,p)$ in terms of $\frac{1}{2}+\varepsilon(d)$, where $\varepsilon(d)$ is the additional term (See Eq. (18), (25), (29), and (42)). This term converges to 0 when $d$ is large. To observe the behaviour, we have plotted $\varepsilon(N)$ (which can be obtained by simply changing the $d$ with $N$ for each symmetric and asymmetric trees) vs $N(d)$.

### v3

@@ A.2 Undirected Binary Trees
$l_{d}^{(r)}=\begin{cases}2d,&\text{if}~r=1\\ 1+3(d-1),&\text{if}~r=2\\ 2+4(d-r+1),&\text{if}~3\leq r\leq d+1\\ 0,&\text{otherwise}.\end{cases}$ (27)
@@
$\left.F^{\text{tel}}_{\text{avg}}(N,p)\right|_{\text{UABT}}$ (30)
$=$ $\frac{1}{2}+\frac{2p(N-1)-(N+3)p^{2}-4p^{3}-(N-1)p^{4}+4p^{\frac{N+3}{2}}+4p^{\frac{N+5}{2}}}{2N(N-1)(1-p)^{2}},$
@@
Here again, $h_{r}$ denotes the number of $r$-length paths starting from the base vertex 1, and at each level, there are two choices for selecting a vertex to include in the path up to depth $r$. So, $h_{r}=2^{r}$. Now, let us denote by $l_{d}^{(r)}$ the number of $r$-length paths in a USBT with depth $d$. In this case, for $r>2d$, there is no $r$-length path as it is an undirected and symmetric tree, and hence $l_{d}^{(r)}=0$ for $r>2d$.
@@
$l_{d}^{(r)}=\begin{cases}3\cdot 2^{d+(\frac{r}{2})-1}-(2d-r+5)\cdot 2^{r-2},&\text{if}~d<r\leq 2d~\text{and}~r~\text{is even}\\ 2^{d+(\frac{r+1}{2})}-(2d-r+5)\cdot 2^{r-2},&\text{if}~d<r\leq 2d~\text{and}~r~\text{is odd}\\ 3\cdot 2^{d+(\frac{r}{2})-1}-(r+3)\cdot 2^{r-2},&\text{if}~r\leq d~\text{and}~r~\text{is even}\\ 2^{d+(\frac{r+1}{2})}-(r+3)\cdot 2^{r-2},&\text{if}~r\leq d~\text{and}~r~\text{is odd}\\ 0,&\text{otherwise}.\end{cases}$ (40)
@@
$\begin{aligned} &\left.F^{\text{tel}}_{\text{avg}}(d,p)\right|_{\text{USBT}}\\ &=\frac{\left[\begin{aligned} &\sum\limits_{\begin{subarray}{c}r\leq d\\ r~\text{even}\end{subarray}}\left(3\cdot 2^{d+\frac{r}{2}-1}-(r+3)\cdot 2^{r-2}\right)\cdot\left(\frac{1+p^{r}}{2}\right)+\sum\limits_{\begin{subarray}{c}r\leq d\\ r~\text{odd}\end{subarray}}\left(2^{d+\frac{r+1}{2}}-(r+3)\cdot 2^{r-2}\right)\cdot\left(\frac{1+p^{r}}{2}\right)+\\ &\sum\limits_{\begin{subarray}{c}d<r\leq 2d\\ r~\text{even}\end{subarray}}\left(3\cdot 2^{d+\frac{r}{2}-1}-(2d-r+5)\cdot 2^{r-2}\right)\cdot\left(\frac{1+p^{r}}{2}\right)+\sum\limits_{\begin{subarray}{c}d<r\leq 2d\\ r~\text{odd}\end{subarray}}\left(2^{d+\frac{r+1}{2}}-(2d-r+5)\cdot 2^{r-2}\right)\cdot\left(\frac{1+p^{r}}{2}\right)\end{aligned}\right]}{\left[\begin{aligned} &\sum\limits_{\begin{subarray}{c}r\leq d\\ r~\text{even}\end{subarray}}\left(3\cdot 2^{d+\frac{r}{2}-1}-(r+3)\cdot 2^{r-2}\right)+\sum\limits_{\begin{subarray}{c}r\leq d\\ r~\text{odd}\end{subarray}}\left(2^{d+\frac{r+1}{2}}-(r+3)\cdot 2^{r-2}\right)+\\ &\sum\limits_{\begin{subarray}{c}d<r\leq 2d\\ r~\text{even}\end{subarray}}\left(3\cdot 2^{d+\frac{r}{2}-1}-(2d-r+5)\cdot 2^{r-2}\right)+\sum\limits_{\begin{subarray}{c}d<r\leq 2d\\ r~\text{odd}\end{subarray}}\left(2^{d+\frac{r+1}{2}}-(2d-r+5)\cdot 2^{r-2}\right)\end{aligned}\right]}\\ &=\frac{1}{2}+\frac{1}{2}\frac{\sum\limits_{\begin{subarray}{c}r=1\\ r~\text{even}\end{subarray}}^{2d}3\cdot 2^{d+\frac{r}{2}-1}\cdot p^{r}+\sum\limits_{\begin{subarray}{c}r=1\\ r~\text{odd}\end{subarray}}^{2d}2^{d+\frac{r+1}{2}}\cdot p^{r}-\sum\limits_{r=1}^{d}(r+3)\cdot 2^{r-2}\cdot p^{r}-\sum\limits_{r=d+1}^{2d}(2d-r+5)\cdot 2^{r-2}\cdot p^{r}}{\sum\limits_{\begin{subarray}{c}r=1\\ r~\text{even}\end{subarray}}^{2d}3\cdot 2^{d+\frac{r}{2}-1}+\sum\limits_{\begin{subarray}{c}r=1\\ r~\text{odd}\end{subarray}}^{2d}2^{d+\frac{r+1}{2}}-\sum\limits_{r=1}^{d}(r+3)\cdot 2^{r-2}-\sum\limits_{r=d+1}^{2d}(2d-r+5)\cdot 2^{r-2}}\\ &=\frac{1}{2}+\frac{1}{2}\frac{\sum\limits_{k=1}^{d}3\cdot 2^{d+k-1}\cdot p^{2k}+\sum\limits_{k=1}^{d}2^{d+k}\cdot p^{2k-1}-\sum\limits_{r=1}^{d}(r+3)\cdot 2^{r-2}\cdot p^{r}-\sum\limits_{t=1}^{d}(d-t+5)\cdot 2^{d+t-2}\cdot p^{d+t}}{\sum\limits_{k=1}^{d}3\cdot 2^{d+k-1}+\sum\limits_{k=1}^{d}2^{d+k}-\sum\limits_{r=1}^{d}(r+3)\cdot 2^{r-2}-\sum\limits_{t=1}^{d}(d-t+5)\cdot 2^{d+t-2}}\\ &=\frac{1}{2}+\frac{1}{2}\frac{\left(3\cdot 2^{d-1}+\frac{2^{d}}{p}\right)\sum\limits_{k=1}^{d}(2p^{2})^{k}-\left(\frac{3}{4}+(d+5)\cdot 2^{d-2}\cdot p^{d}\right)\sum\limits_{k=1}^{d}(2p)^{k}+\left(2^{d-2}\cdot p^{d}-\frac{1}{4}\right)\sum\limits_{k=1}^{d}k(2p)^{k}}{\left(3\cdot 2^{d-1}+2^{d}\right)\sum\limits_{k=1}^{d}2^{k}-\left(\frac{3}{4}+(d+5)\cdot 2^{d-2}\right)\sum\limits_{k=1}^{d}2^{k}+\left(2^{d-2}-\frac{1}{4}\right)\sum\limits_{k=1}^{d}k2^{k}}\\ &=\begin{cases}\displaystyle\frac{1}{2}+\frac{14(2^{d}-1)-d(d+8)}{8(2^{d}-1)(2^{d+1}-1)},&\text{if }~p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{d\cdot 2^{d+1}-\sqrt{2}(2^{d/2}-1)\left[(\sqrt{2}-1)\left\{3+5\cdot 2^{d/2}\right\}+(2^{d/2}-1)\right]}{8(3-2\sqrt{2})(2^{d}-1)(2^{d+1}-1)},&\text{if }~p=\frac{1}{\sqrt{2}},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{2^{d-1}p(3p+2)\left((2p^{2})^{d}-1\right)}{(2p^{2}-1)(2^{d}-1)(2^{d+1}-1)}-\frac{p\left((2p)^{d}-1\right)\left[(2p-1)\left\{3+5\cdot(2p)^{d}\right\}+\left((2p)^{d}-1\right)\right]}{4(2p-1)^{2}(2^{d}-1)(2^{d+1}-1)},&\text{otherwise}.\end{cases}\end{aligned}$
Therefore, we can write,
$\left.F^{\text{tel}}_{\text{avg}}(d,p)\right|_{\text{USBT}}=\frac{1}{2}+\varepsilon(d).$ (41)
@@
We did not consider the case $p=1$ while computing the average fidelity in the above four trees, namely DABT, DSBT, UABT and USBT, since then we would have $\mathcal{F}_{r}=\frac{1+p^{r}}{2}=1$, hence
$F^{\text{tel}}_{\text{avg}}=\frac{\sum\limits_{r}l_{d}^{(r)}\mathcal{F}_{r}}{\sum\limits_{r}l_{d}^{(r)}}=1.$
@@ Appendix B Large $N$ limit
To understand the large $N$ behaviour for all the cases of binary trees, we consider the formula of $F_{\mathrm{avg}}^{\mathrm{tel}}$. For any value of $p<1$, we can write the formula of $F_{\mathrm{avg}}^{\mathrm{tel}}(d,p)$ in terms of $\frac{1}{2}+\varepsilon(d)$, where $\varepsilon(d)$ is the additional term (See Eq. (18), (25), (29), and (41)). This term converges to 0 when $d$ is large. To observe the behaviour, we have plotted $\varepsilon(N)$ (which can be obtained by simply changing the $d$ with $N$ for each symmetric and asymmetric trees) vs $N(d)$.